Please Help. Thanks.
Find the y-intercept for -3x – 2y = 12

Answer by varralovesjbieber(2)

-3x-2y=12
-3x-2(0)=12
-3x=12
-3x/-3=12/-3
x=-4
-3(0)-2y=12
-2y=12
-2y/-2=12/-2
y=-6